# Engineering a Josephson junction chain for the simulation of the clock model

### v1

@@ V.3 Chiral model at $n_{g}=0$ and the floating phase
Figure 5: Summary of the phase diagram in the $E_{c}-A$ plane for ${J_{0}}=5E_{J}$: (a) phase winding $\mathcal{W}$, (b) order parameter, (c) current density, (d) central charge. The continuous black line marks the analytical estimate of the phase transitions at small charging energies, where the charge truncation would induce nonphysical effects in our simulations. The colored arrows indicate the horizontal cuts of the phase diagram used for Figs. 6 and 8. The system size is $L=100$.
@@
The green curve in Fig. 6(a) shows that the paramagnetic phase immediately acquires a non-commensurate winding $k$ for $A\neq 0$. This winding converges to $A$ sufficiently deep in the gapped disordered phase. When considering the evolution of the ferromagnetic phase as a function of $A$, there are two distinct cases that are separated by a limiting value of the charging energy $\tilde{E}_{C}<E^{*}_{C}$.
@@
Concerning the transition from the floating phase to the paramagnetic phase, we observe that the winding $k$ does not display any discontinuity. Based on the field theory description (10) this transition happens when the sine-Gordon potential $\cos(2\theta)$ becomes relevant ($K=2$). Since this potential is not affected by the incommensurability parameter $A$, we conclude that this transition is of the Berezinskii-Kosterlitz-Thouless (BKT) kind. Therefore, when considering a vertical cut of the phase diagrams 5 and 11 such that all three phases are present, the two gapped/gapless phase transitions are of different nature: the first, from an ordered to the floating phase, is a non-conformal commensurate-incommensurate transition; the second, from the floating to the paramagnetic phase, is instead of the BKT class. This situation is analogous to the study of Rydberg atom systems with $\mathbb{Z}_{3}$ symmetry [12]. The BKT nature of the second transition, however, is not stable under the introduction of the other incommensurability parameter $n_{g}$, which turns also this phase transition into the commensurate/incommensurate kind.
@@
We show the expectation value of $I_{j}$ averaged over the position in Fig. 8 which, as expected, is periodic in $A$ with a period $32\pi/$ 3. The ordered phases are characterized by a Meissner current linear in $A$, whose appearance is made possible by the triple Josephson junctions connected with a superconducting background. The commensurate-incommensurate phase transitions into the gapless phase are signaled by a sharp cusp in the current dependence on $A$, analogous to bosonic ladder models [63, 64]. The floating phase corresponds indeed to an incommensurate vortex phase and its average current per link decreases to zero for $A$ approaching $\pi/3$. Finally the paramagnetic phase corresponds to a Mott insulator for the Cooper pairs. In this phase the current is exponentially suppressed with the system size and shows little or no dependence on $A$.
@@ VI The effect of the induced charge
This implies that, for $n_{g}\neq 0$, the transition from ordered superconducting and disordered insulating phases driven by $E_{c}$ is no longer of the Potts kind, analogously to the case with generic $A$. Moreover, for generic $A$ and $n_{g}$, the transitions from the gapless floating phase to the gapped phases is always of the commensurate/incommensurate kind, either for the charge degrees of freedom (thus the $\theta$ field) or for the phase $\varphi$ degrees of freedom.
@@ VII Perturbations and disorder
To provide rigorous modeling of the experimental platforms for the observation of these phases of matter, one needs to account also for the unavoidable disorder that affects the parameters of the JJAs, despite the accuracy of modern fabrication techniques.
@@
Disorder may have a major influence when approaching the Potts critical point in the limit $n_{g}=A=0$. In this scenario, we must distinguish two kinds of perturbations: perturbations that leave the system invariant under the $\mathbb{Z}_{3}$ symmetry, such as disorder in the $E_{J},{J_{0}},E_{c},A$ and $n_{g}$ parameters, may lead to relatively weak deformations of the low-energy scaling limit of the model since they are coupled to operators with larger scaling dimensions based on the scaling analysis of disordered systems [69]. Specifically, quenched disorder in the parameters $M$ and ${J_{0}}$ in Eq. (10) can lead to different scaling exponents similarly to a random transverse field for the quantum Ising model [70]. On the other hand, perturbations that break the $\mathbb{Z}_{3}$ symmetry, instead, are expected to have an effect similar to the detuning of the single island potential we studied in Sec. VII. In this respect, disorder is expected to introduce an additional length scale in the model, which diverges in the clean systems. The Potts critical behavior can be observed only if this length scale is sufficiently large to be comparable with the system size. A complete understanding of the role of disorder at the Potts transition requires a deep analysis in terms of renormalization group and CFTs.
@@
The role of disorder becomes even more complex when considering the commensurate/incommensurate phase transitions and the floating phase. For instance, in the regime of low $E_{c}$ at $A\sim\pi/3$, we expect that charge disorder is strongly suppressed, and the incommensurability introduced by $A$ additionally suppresses the disorder in the onsite potential. In this case, it may even be possible that a further disorder in ${J_{0}}$ or in the parameter $A$ could extend the floating phase. Concerning the Coulomb-dominated regime at large $E_{c}$, instead, the role of charge disorder becomes important; its effect can be understood based on the mapping to the Bose-Hubbard model and can lead to the formation of additional Bose glass phases [71] (see also the numerical results in Ref. [72]).
@@
As recent experiments in long JJAs [34] estimated a disorder strength in the Josephson couplings below 10%, we are confident that precise fabrication of the device allows for the observation of the desired phenomenology.
@@ VIII Conclusions
These superconducting systems provide a solid-state architecture for quantum simulations that combines two key features: (i) their dynamics is determined by two controllable incommensurability parameters, which correspond to the induced charges in the superconducting islands and the magnetic fluxes in their geometry; (ii) they realize a global discrete $\mathbb{Z}_{3}$ symmetry which, differently from Rydberg atom experiments [8, 9] and triangular lattice antiferromagnets [73], is not related to spacial symmetries as translational or rotational invariance. The $\mathbb{Z}_{3}$ symmetric and symmetry-broken states of these Josephson junction chains are indeed associated with different configurations of their superconducting phases. This symmetry can be engineered already at the level of a single Josephson junction element by taking advantage of the gate-tunability of the energy-phase relation recently demonstrated in hybrid parallel junctions [45].
@@
In a broader context, hybrid JJAs are compelling elements for pushing superconducting quantum circuits beyond the qubit realm. We observe, in particular, that the triple Josephson junction can be operated as a qutrit, if suitably coupled with a microwave resonator as for standard transmons. In this scenario, the computational states are encoded in the minima of the Josephson potential as in parity [74] and modern flux qubits [75]. Single qutrit-gates could be straightly generalized from their qubit counterparts [75, 76], while entangling gates and readout methods would require a more careful investigation.
@@ Acknowledgements.
We warmly thank E. Cobanera, G. Ortiz, M. Rizzi, and N. Tausendpfund for useful discussions. MW has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work was supported by Q@TN, the joint lab between the University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. M.B. was supported by the Villum foundation (Research Grant No. 25310).
@@ Appendix A Semiclassical estimate of the phase slips in the triple junction
We remark that, in our mapping to the quantum clock model, we neglected the simultaneous phase slips of neighboring islands, which can be estimated by extending the results in Ref. [77] and would result in transition amplitudes with an exponential suppression in ${J_{0}}/E_{c}$ stronger than Eq. (32). Our numerical MPS simulations, however, do not rely on any of these semiclassical simulations, and capture the ground states of the system independently on the ratio ${J_{0}}/E_{c}$.
@@ Appendix B Kink mass in the effective field theory description
Replacing the perturbative expression (11) for the Luttinger liquid parameters yelds the mass of the kink $M_{\rm kink}$ adopted in the main text, from which we derive the parameter $M$ in Eq. (12).
@@ Appendix D Perturbations of the single-island potential
In Sec. VII we discussed the main effect of the perturbations $\delta J_{i}$ in the potential (29) in terms of a splitting in energy of the semiclassical potential minima that corresponds to the longitudinal field in Eq. (30).
@@
Considering the unperturbed system at the Potts phase transition, the perturbations $\delta h_{\tau}$ and $\delta h_{\sigma\tau}$ are relevant but subleading with respect to the longitudinal field term with coefficient $\delta h_{\sigma}$. In general we expect a scaling of the order parameter away from the critical point dictated by the universal behavior:
@@
where $a$ runs over the relevant primary perturbations determined by the Potts conformal field theory, $D_{a}$ are the scaling dimensions of the corresponding primary fields and $A_{a}$ are non-universal constants. When considering the mapping of the clock model into the Potts conformal field theory (see Ref. [26] for an advanced analysis), the perturbation in $\delta h_{\sigma}$ matches the primary field with dimension $D_{\sigma}=\frac{2}{15}$ and the displacement $\delta h_{\tau}$ corresponds to the primary field with dimension $D_{\tau}=4/5$. The perturbation $\delta h_{\sigma\tau}$, instead, is not directly related to a primary operator. Its dominant term should however be characterized by the same scaling of the longitudinal field.

### v2

@@ V.3 Chiral model at $n_{g}=0$ and the floating phase
Figure 5: Summary of the phase diagram in the $E_{c}-A$ plane for ${J_{0}}=5E_{J}$: (a) phase winding $k$, (b) order parameter, (c) current density, (d) central charge. The continuous black lines mark the analytical estimate of the phase transitions at small charging energies, where the charge truncation would induce nonphysical effects in our MPS simulations. The colored arrows indicate the horizontal cuts of the phase diagram used for Figs. 6 and 8. The system size is $L=100$.
@@
The green curve in Fig. 6(a) shows that the paramagnetic phase immediately acquires a non-commensurate winding $k$ for $A\neq 0$. This winding converges to $A$ sufficiently deep in the gapped disordered phase. When considering the evolution of the ferromagnetic phase as a function of $A$, there are two distinct cases that are separated by a limiting value of the charging energy $\tilde{E}_{c}<E_{c}^{*}$.
@@
Concerning the transition from the floating phase to the paramagnetic phase, we observe that the winding $k$ does not display any discontinuity. Based on the field theory description (10) this transition happens when the sine-Gordon potential $\cos(2\theta)$ becomes relevant ($K=2$). Since this potential is not affected by the incommensurability parameter $A$, we conclude that this transition is of the Berezinskii-Kosterlitz-Thouless (BKT) kind. Therefore, when considering a vertical cut of the phase diagrams 5 and 11 such that all three phases are present, the two gapped/gapless phase transitions are of different nature: the first, from an ordered to the floating phase, is a non-conformal commensurate-incommensurate transition; the second, from the floating to the paramagnetic phase, is instead of the BKT class. This situation is analogous to the study of Rydberg atom systems with $\mathbb{Z}_{3}$ symmetry [12]. The BKT nature of the second transition, however, is not stable under the introduction of the other incommensurability parameter $n_{g}$, which turns also this phase transition into the commensurate-incommensurate kind.
@@
We show the expectation value of $I_{j}$ averaged over the position in Fig. 8 which, as expected, is periodic in $A$ with a period $32\pi/$ 3. The ordered phases are characterized by a Meissner current linear in $A$, whose appearance is made possible by the triple Josephson junctions connected with a superconducting background. The commensurate-incommensurate phase transitions into the gapless phase are signaled by a sharp cusp in the current dependence on $A$, analogous to bosonic ladder models [68, 69]. The floating phase corresponds indeed to an incommensurate vortex phase and its average current per link decreases to zero for $A$ approaching $\pi/3$. Finally, the paramagnetic phase corresponds to a Mott insulator for the Cooper pairs. In this phase, the current is exponentially suppressed with the system size and shows little or no dependence on $A$.
@@ VI The effect of the induced charge
This implies that, for $n_{g}\neq 0$, the transition from ordered superconducting and disordered insulating phases driven by $E_{c}$ is no longer of the Potts kind, analogously to the case with generic $A$. Moreover, for generic $A$ and $n_{g}$, the transitions from the gapless floating phase to the gapped phases is always of the commensurate-incommensurate kind, either for the charge degrees of freedom (thus the $\theta$ field) or for the phase $\varphi$ degrees of freedom.
@@ VII Perturbations and disorder
Another perturbation of the Hamiltonian (7) that may affect the experimental platforms is the onset of repulsive interactions between the islands. As mentioned in Sec. III, these interactions are dictated by the inverse capacitance matrix of the system $C^{-1}$. Among them, the most relevant are given by the nearest-neighbor terms of the form $E_{cc}\left(N_{j}-n_{g}\right)\left(N_{j+1}-n_{g}\right)$ characterized by an energy scale $E_{cc}<E_{c}/2$ (in particular, for hybrid systems, $E_{cc}<E_{c}/4$ in the realistic estimates presented in Ref. [52]). These mutual electrostatic interactions do not lead to qualitative differences in the physics of the Josephson junction arrays for systems that are dominated by the Josephson potentials. In particular, nearest-neighbor interactions can be neglected in the transmon-like regime discussed in Sec. II as the related two-island phase slips ($\tau_{j}\tau_{j+1}$) display amplitudes that are exponentially suppressed approximately by the factor $\exp\left[{-\sqrt{\frac{256J_{0}\cos(3A/2)}{9(E_{c}+2E_{cc})}}}\right]$, to be compared to $h_{\tau}$ in Eq. (6) (see Appendix A). We observe, however, that the mutual repulsive interactions lower the value of the Luttinger parameter $K$, thus lowering the critical value $E_{c}^{*}$ and shrinking the extension of the incommensurate phase appearing around $A=\pi/3$ for low $E_{c}$. Nearest-neighbor interactions, instead, may play a more decisive role in the Coulomb-dominated regimes. In particular they favor the onset of staggered charge density wave for $n_{g}\sim 1/2$ in the Bose-Hubbard model of Cooper pairs emerging for $E_{c}\gg E_{J},J_{0}$ [53]. Analogously to the Mott insulating states, these staggered states suppress the charge fluctuations and constitute insulating phases that compete with the gapless superfluid phase described in Sec. VI.
@@
Finally, to provide rigorous modeling of the experimental platforms for the observation of these phases of matter, one needs to account also for the unavoidable disorder that affects the parameters of the JJAs, despite the accuracy of modern fabrication techniques.
@@
Disorder may have a major influence when approaching the Potts critical point in the limit $n_{g}=A=0$. In this scenario, we must distinguish two kinds of perturbations: perturbations that leave the system invariant under the $\mathbb{Z}_{3}$ symmetry, such as disorder in the $E_{J},{J_{0}},E_{c},A$ and $n_{g}$ parameters, may lead to relatively weak deformations of the low-energy scaling limit of the model since they are coupled to operators with larger scaling dimensions based on the scaling analysis of disordered systems [73]. Specifically, quenched disorder in the parameters $M$ and ${J_{0}}$ in Eq. (10) can lead to different scaling exponents similarly to a random transverse field for the quantum Ising model [74]. On the other hand, perturbations that break the $\mathbb{Z}_{3}$ symmetry, instead, are expected to have an effect similar to the detuning of the single island potential $\delta J_{i}$ discussed above. In this respect, disorder is expected to introduce an additional length scale in the model, which diverges in the clean systems. The Potts critical behavior can be observed only if this length scale is sufficiently large to be comparable with the system size. A complete understanding of the role of disorder at the Potts transition, however, requires a deep analysis in terms of renormalization group and CFTs and goes beyond the scope of this work.
@@
The role of disorder becomes even more complex when considering the commensurate-incommensurate phase transitions and the floating phase. For instance, in the regime of low $E_{c}$ at $A\sim\pi/3$, we expect that charge disorder is strongly suppressed, and the incommensurability introduced by $A$ additionally suppresses the disorder in the onsite Josephson potential. In this case, it may even be possible that a further disorder in ${J_{0}}$ or in the parameter $A$ could extend the floating phase. Concerning the Coulomb-dominated regime at large $E_{c}$, instead, the role of charge disorder becomes important; its effect can be understood based on the mapping to the Bose-Hubbard model and can lead to the formation of additional Bose glass phases [75] (see also the numerical results in Ref. [76]).
@@
As recent experiments in long JJAs [34] estimated a disorder strength in the Josephson couplings below 10%, however, we are confident that a state-of-the-art fabrication of the device allows for the observation of all the main features of the phases of matter discussed in the previous sections.
@@ VIII Conclusions
These superconducting systems provide a solid-state architecture for quantum simulations that combines two key features: (i) their dynamics is determined by two controllable incommensurability parameters, which correspond to the induced charges in the superconducting islands and the magnetic fluxes in their geometry; (ii) they realize a global discrete $\mathbb{Z}_{3}$ symmetry which, differently from Rydberg atom experiments [8, 9] and triangular lattice antiferromagnets [77], is not related to spacial symmetries as translational or rotational invariance. The $\mathbb{Z}_{3}$ symmetric and symmetry-broken states of these Josephson junction chains are indeed associated with different configurations of their superconducting phases. This symmetry can be engineered already at the level of a single Josephson junction element by taking advantage of the gate-tunability of the energy-phase relation recently demonstrated in hybrid parallel junctions and loops[46, 49].
@@
In a broader context, hybrid JJAs are compelling elements for pushing superconducting quantum circuits beyond the qubit realm. We observe, in particular, that the triple Josephson junction we described can be operated as a qutrit, if suitably coupled with a microwave resonator as for standard transmons. In this scenario, the computational states are encoded in the minima of the Josephson potential as in parity [78] and modern flux qubits [79]. Single qutrit-gates could be straightly generalized from their qubit counterparts [79, 80], while entangling gates and readout methods would require a more careful investigation. In this respect, the superconducting arrays that we discussed constitute a starting point for the realization of tunable solid-state qutrit architectures.
@@ Acknowledgements.
We warmly thank E. Cobanera, L. Banszerus, G. Ortiz, M. Rizzi, N. Tausendpfund, and S. Vaitiekenas for useful discussions. MW has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101080086 NeQST. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work was supported by Q@TN, the joint lab between the University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. M.B. and L.M. have been supported by the Villum Foundation (Research Grant No. 25310). L.M. has been supported by the research grant “PARD 2023”, and “Progetto di Eccellenza 23-27” funded by the Department of Physics and Astronomy G. Galilei, University of Padua.
Data Availability
The data presented in this article is available from [81].
@@ Appendix A Semiclassical estimate of the phase slips in the triple junction
In our mapping to the quantum clock model, we neglected the simultaneous phase slips of neighboring islands. Analogously to the physics of Josephson junction chains hosting $\mathbb{Z}_{N}$ parafermions [17, 24], these two-island phase slips would amount to operators of the form $\tau_{j}\tau_{j+1}$ in the clock model limit. They arise, in particular, when neighboring islands are interacting through an electrostatic repulsion of the form $E_{cc}\left(N_{j}-n_{g}\right)\left(N_{j+1}-n_{g}\right)$.
The amplitude of these two-island phase slips can be estimated by adopting the approach in Refs. [82, 83]. In particular, it can be easily evaluated in the limit $E_{J}>J_{0}\gg E_{c}>E_{cc}$. In this scenario, we can consider two neighboring islands as a single superconducting element sharing the same phase. Such element acquires an effective charging energy $\tilde{E}_{c}=\frac{2E_{c}+E_{cc}}{4}<E_{c}$ and is connected to the superconducting background through an effective SQUID for triplets of Cooper pairs with potential $-\tilde{J}_{0}\cos 3\varphi$, with $\tilde{J}_{0}=2J_{0}\cos(3A/2)$
By repeating the previous calculations, we derive that the amplitude of the two-island phase slips scales as $\tilde{J}_{0}^{3/4}\tilde{E}_{c}^{1/3}\mathrm{e}^{-\sqrt{\frac{32\tilde{J}_{0}}{9\tilde{E}_{c}}}}$. For small $A$, this amplitude is strongly suppressed with respect to $h_{\tau}$ in Eq. (32), thus bearing only minor corrections to the Potts phase transition.
We emphasize that our numerical MPS simulations do not rely on any of these semiclassical approximations, and capture the ground states of the system independently on the ratio ${J_{0}}/E_{c}$.
@@ Appendix B Kink mass in the effective field theory description
Replacing the perturbative expression (11) for the Luttinger liquid parameters yields the mass of the kink $M_{\rm kink}$ adopted in the main text, from which we derive the parameter $M$ in Eq. (12).
@@ Appendix D Perturbations of the single-island potential
In Sec. VII we discussed the main effect of the perturbations $\delta J_{i}$ in the potential (29) in terms of a splitting in energy of the semiclassical potential minima that correspond to the longitudinal field in Eq. (30).
@@
Considering the unperturbed system at the Potts phase transition, the perturbations $\delta h_{\tau}$ and $\delta h_{\sigma\tau}$ are relevant but subleading with respect to the longitudinal field term with coefficient $\delta h_{\sigma}$. In general, we expect a scaling of the order parameter away from the critical point dictated by the universal behavior:
@@
where $a$ runs over the relevant primary perturbations determined by the Potts conformal field theory, $D_{a}$ are the scaling dimensions of the corresponding primary fields and $A_{a}$ are non-universal constants. When considering the mapping of the clock model into the Potts conformal field theory (see Ref. [26] for advanced analysis), the perturbation in $\delta h_{\sigma}$ matches the primary field with dimension $D_{\sigma}=\frac{2}{15}$ and the displacement $\delta h_{\tau}$ corresponds to the primary field with dimension $D_{\tau}=4/5$. The perturbation $\delta h_{\sigma\tau}$, instead, is not directly related to a primary operator. Its dominant term should however be characterized by the same scaling of the longitudinal field.